# Zeroth-order log-concave sampling

### v1

@@ Abstract
We study the zeroth-order query complexity of log-concave sampling, specifically uniform sampling from convex bodies using membership oracles. We propose a simple variant of the proximal sampler that achieves the query complexity with matched Rényi orders between the initial warmness and output guarantee. Specifically, for any $\varepsilon>0$ and $q\geq 2$, the sampler, initialized at $\pi_{0}$, outputs a sample whose law is $\varepsilon$-close in $q$-Rényi divergence to $\pi$, the uniform distribution over a convex body in $\mathbb{R}^{d}$, using $\widetilde{O}(qM_{q}^{q/(q-1)}d^{2}\,\lVert\operatorname{cov}\pi\rVert\log\frac{1}{\varepsilon})$ membership queries, where $M_{q}=\lVert\text{d}\pi_{0}/\text{d}\pi\rVert_{L^{q}(\pi)}$.
We further introduce a simple annealing scheme that produces a warm start in $q$-Rényi divergence (i.e., $M_{q}=O(1)$) using $\widetilde{O}(qd^{2}R^{3/2}\,\lVert\operatorname{cov}\pi\rVert^{1/4})$ queries, where $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$. This interpolates between known complexities for warm-start generation in total variation and Rényi-infinity divergence. To relay a Rényi warmness across the annealing scheme, we establish hypercontractivity under simultaneous heat flow and translate it into an improved mixing guarantee for the proximal sampler under a logarithmic Sobolev inequality. These results extend naturally to general log-concave distributions accessible via evaluation oracles, incurring additional quadratic queries.
@@ 1 Introduction
We study the zeroth-order query complexity of sampling from an arbitrary log-concave distribution $\pi$, assuming access to an evaluation oracle for $-\log\pi$ (up to an additive constant). This is a very general setting for log-concave sampling—for example, it includes uniform sampling from convex bodies, which serves as a running example throughout this paper. Results in this general setting provide a theoretical baseline, highlighting the benefits of additional assumptions often made in other sampling work, such as log-smoothness (i.e., $-\nabla\log\pi$ is Lipschitz) or access to first-order oracles (i.e., $-\nabla\log\pi$). Indeed, much of our work is motivated by gaps between this general case and more favorable settings where first-order oracles are available for positive and log-smooth distributions. In practical terms, zeroth-order samplers are natural default algorithms, since evaluation oracles typically require only function value computations, making them the easiest to implement.
@@
We now make the problem setup more explicit, focusing for now on uniform sampling. Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body (i.e., a convex and compact set with non-empty interior) containing the $d$-dimensional unit ball $B_{1}$. Given access to the membership oracle for $\mathcal{K}$ (which answers whether a queried point belongs to $\mathcal{K}$) [GLS93], we ask: how can we sample uniformly from $\mathcal{K}$, and how many oracle queries are required? More formally, given a target accuracy $\varepsilon>0$ and a probability divergence (or distance) $\mathsf{D}$—such as the total variation ($\mathsf{TV}$) distance or $\chi^{2}$-divergence—what is the query complexity of generating a sample whose law is $\varepsilon$-close to the uniform distribution $\pi$ over $\mathcal{K}$ in $\mathsf{D}$?
@@
In this work, we propose a simple uniform sampler that essentially subsumes prior methods developed over the past three decades. Its simplicity allows us to establish the first balanced complexity guarantee in a general setting, meaning that the Rényi orders of the initial warmness and the output metric are the same. This result parallels guarantees found in recent sampling research in the log-smooth setting. Next, we provide a simple annealing scheme that, for the first time, generates a warm start in $q$-Rényi divergence. To this end, we prove hypercontractivity under simultaneous heat flow, given a logarithmic Sobolev inequality. Finally, we extend these results to log-concave distributions under evaluation oracles. Alongside, we present a streamlined analysis leveraging powerful tools from diverse mathematical fields.
1.1 Backgrounds and results
@@
In this section, $\pi_{\mathcal{K}}$ denotes the uniform distribution over a convex body $\mathcal{K}$ with diameter $D$, and $\pi$ refers to log-concave distributions over $\mathbb{R}^{d}$. For clarity, we recall the $q$-Rényi divergence, defined as $\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{q}{q-1}\log\,\lVert\nicefrac{{\mathrm{d}\mu}}{{\mathrm{d}\nu}}\rVert_{L^{q}(\nu)}$ for $q>1$, and its monotonicity in $q$, and $2\,\lVert\cdot\rVert_{\mathsf{TV}}^{2}\leq\mathsf{KL}=\lim_{q\downarrow 1}\mathcal{R}_{q}\leq\mathcal{R}_{2}=\log(1+\chi^{2})\leq\chi^{2}\leq\mathcal{R}_{\infty}=\log\operatorname{ess\,sup}_{\nu}\nicefrac{{\mathrm{d}\mu}}{{\mathrm{d}\nu}}$. See §1.4 for more details. When a sampler is initialized from an initial distribution $\nu_{0}$ toward a target distribution $\nu$, we define the warmness parameter as $M_{q}=\lVert\mathrm{d}\nu_{0}/\mathrm{d}\nu\rVert_{L^{q}(\nu)}$.
@@ 1.1.1 Uniform sampling from a warm start
The $\mathsf{Ball\ walk}$ and $\mathsf{Hit\text{-}and\text{-}Run}$ are well-studied random walks (i.e., Markov chain Monte Carlo methods) for uniform sampling. These algorithms are analyzed through the conductance framework, where a lower bound on conductance yields a mixing-time guarantee. Establishing such a lower bound involves two ingredients: one-step coupling (i.e., the closeness of transition kernels at two nearby points) and a Cheeger isoperimetric inequality for the target distribution. The Cheeger constant of a probability measure $\pi$ over $\mathbb{R}^{d}$ is defined as
@@
The $\mathsf{Ball\ walk}$ was introduced by Lovász [Lov90]; in each step, it samples a random point uniformly from a ball of fixed radius around the current point, accepting it if the point lies within $\mathcal{K}$. Later, Lovász and Simonovits established that the $\mathsf{Ball\ walk}$ satisfies $\mathcal{R}_{\infty}\to\mathsf{TV}$ with complexity $\operatorname{poly}(d,D,M_{\infty}/\varepsilon)$ [LS93, Remark on p. 398], meaning the initial warmness should be given in $\mathcal{R}_{\infty}$-divergence and the final output is $\varepsilon$-close to $\pi_{\mathcal{K}}$ in $\mathsf{TV}$-distance. They also analyzed an indirect approach that uses the $\mathsf{Ball\ walk}$ (with a Metropolis filter) to sample from a log-concave distribution and then post-processes the sample to bring it closer to $\pi_{\mathcal{K}}$. The resulting algorithm, also of type $\mathcal{R}_{\infty}\to\mathsf{TV}$, was shown to have query complexity $\operatorname{poly}(d,D,\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$.
@@
This approach was further improved by Kannan, Lovász, and Simonovits [KLS97], who introduced the $\mathsf{Speedy\ walk}:\mathcal{R}_{\infty}\to\mathsf{TV}$. Similar to [LS93], it begins by sampling from a biased log-concave variant $\pi^{\prime}$ (of $\pi_{\mathcal{K}}$), which is invariant under the following per-step transition: the next point $x_{i+1}$ is sampled uniformly from $\mathcal{K}\cap B_{\delta}(x_{i})$ via rejection sampling, using the uniform distribution over $B_{\delta}(x_{i})$ as the proposal. Then, a post-processing step is used to reduce the bias. They showed that the mixing time of this chain is $O(d^{2}C_{\mathsf{Ch}}^{-2}(\pi^{\prime})\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$ and that each rejection sampling step requires $O(M_{\infty})$ queries, leading to an overall complexity of $\widetilde{O}(M_{\infty}d^{2}D^{2}\log\nicefrac{{1}}{{\varepsilon}})$ (due to $C_{\mathsf{Ch}}(\pi)\gtrsim D^{-1}$ [LS93]). Hereafter, by following convention, the $\mathsf{Ball\ walk}$ refers to the $\mathsf{Speedy\ walk}$ with the post-processing step. Later, progress on the Kannan-Lovász-Simonovits (KLS) conjecture [KLS95] improved this bound to $C_{\mathsf{Ch}}^{2}(\pi)\gtrsim(\lVert\operatorname{cov}\pi\rVert\log d)^{-1}$ [Kla23]. As a result, the complexity of the $\mathsf{Ball\ walk}$ is now $\widetilde{O}(M_{\infty}d^{2}\lVert\operatorname{cov}\pi\rVert\log\nicefrac{{1}}{{\varepsilon}})$.
@@
The $\mathsf{Hit\text{-}and\text{-}Run}$ algorithm [Smi84] was first analyzed by Lovász [Lov99] via $s$-conductance [LS90], with subsequent refinements by Lovász and Vempala [LV06b]. Each step selects a random chord passing the current point and moves to a random point on the chord. Here, uniform sampling from the chord is implemented via rejection sampling with complexity $O(\log D)$, where the proposal distribution is uniform over a slightly longer chord (found by binary search). They established a novel Cheeger isoperimetric inequality in terms of the cross-cut ratio (rather than Euclidean distance) and showed $\mathsf{Hit\text{-}and\text{-}Run}:\mathcal{R}_{2}\to\mathcal{R}_{2}$ with complexity $\widetilde{O}(d^{2}D^{2}\log\nicefrac{{M_{2}}}{{\varepsilon}})$, making it the first random walk with a balanced guarantee. Thus, from an $O(1)$-warm start in $\mathcal{R}_{\infty}(\geq\mathcal{R}_{2})$, the $\mathsf{Ball\ walk}$ achieves faster mixing (and better query complexity) than $\mathsf{Hit\text{-}and\text{-}Run}$.
@@
However, despite these advances, a complete and clean understanding of this fundamental problem is still lacking. First, the $\mathsf{Ball\ walk}$ takes an indirect approach to uniform sampling: as noted earlier, it samples from a biased distribution to obtain a mixing guarantee with $\log\nicefrac{{M_{\infty}}}{{\varepsilon}}$-dependence, followed by a post-processing step to reduce the bias. Second, conductance-based analyses require an algorithmic modification making the chains lazy (i.e., the chain remains at its current point with some probability). To achieve greater simplicity and elegance, it would be desirable to develop a direct sampling algorithm that avoids these additional processing steps. Finally, the highly unbalanced guarantee of the $\mathsf{Ball\ walk}$ (i.e., $\mathcal{R}_{\infty}\to\mathsf{TV}$) appears unnatural. As with $\mathsf{Hit\text{-}and\text{-}Run}$, we are led to ask whether we can obtain a balanced guarantee using the improved Cheeger constant $C_{\mathsf{Ch}}^{2}(\pi)\gtrsim(\lVert\operatorname{cov}\pi\rVert\log d)^{-1}$. This question becomes even more natural when considering recent developments in sampling under log-smoothness.
@@
Modern research in sampling has made substantial progress by analyzing discretizations of continuous-time processes. A central example is the Langevin diffusion $(X_{t})_{t\geq 0}$ in $\mathbb{R}^{d}$, described by the stochastic differential equation
@@
and the smallest such $C$ is called the Poincaré constant $C_{\mathsf{PI}}(\pi)$. Furthermore, the seminal work of Jordan, Kinderlehrer, and Otto [JKO98] interpreted Langevin diffusion as the $2$-Wasserstein gradient flow of $\mathsf{KL}(\cdot\mathbin{\|}\nu)$, revealing a connection to the logarithmic Sobolev inequality, which is strictly stronger than the Poincaré inequality. A probability measure $\pi$ on $\mathbb{R}^{d}$ is said to satisfy a logarithmic Sobolev inequality with constant $C$ if for any locally Lipschitz function $f\in L^{2}(\pi)$,
@@
and the smallest such $C$ is referred to as the log-Sobolev constant $C_{\mathsf{LSI}}(\pi)$. Under ($\mathsf{PI}$) and ($\mathsf{LSI}$) respectively, the Langevin process satisfies the exponential convergences:
$\chi^{2}(\nu_{t}\mathbin{\|}\nu)\leq\exp\bigl{(}-\frac{2t}{C_{\mathsf{PI}}(\nu)}\bigr{)}\,\chi^{2}(\nu_{0}\mathbin{\|}\nu)\,,\quad\text{and}\quad\mathsf{KL}(\nu_{t}\mathbin{\|}\nu)\leq\exp\bigl{(}-\frac{2t}{C_{\mathsf{LSI}}(\nu)}\bigr{)}\,\mathsf{KL}(\nu_{0}\mathbin{\|}\nu)\,,$ (1.1)
which were later extended to $\mathcal{R}_{q}$-divergence in [CLL19, VW23].
With bounded smoothness (i.e., $\beta:=\lVert\nabla^{2}V\rVert<\infty$), a short-time discretization of the Langevin process yields an algorithm that preserves the balanced guarantees of the continuous process. Specifically, it has been shown to converge in $\mathsf{KL}$ (i.e., $\mathcal{R}_{1}\to\mathcal{R}_{1}$) with complexity $\operatorname{poly}(d/\varepsilon,\beta,C_{\mathsf{LSI}}(\nu),\log\mathsf{KL}(\nu_{0}\mathbin{\|}\nu))$ [VW23]. This result was further extended to $\mathcal{R}_{q}$-divergence [CEL+24] achieving a query complexity of $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{PI}}(\nu),\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$ or $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{LSI}}(\nu),\log\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$. We note that the dependence on $\nicefrac{{1}}{{\varepsilon}}$ can be made polylogarithmic by incorporating a Metropolis filter (e.g., $\mathsf{MALA}$ [CLA+21, WSC22, CG23a]) or by considering $\mathsf{MHMC}$ [CV22, CDWY20, CG23b]). For a comprehensive review, we refer readers to the monograph [Che25].
The complexity of the $\mathsf{Ball\ walk}$ can be partially described as $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{PI}}(\nu),\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$. In fact, for log-concave probability measures $\pi$, it holds that
@@
Rise of the proximal sampler.
Following this line of work, the $\mathsf{Proximal\ sampler}$ ($\mathsf{PS}$) [LST21] has brought substantial progress to sampling research. For a target distribution $\pi^{X}\propto e^{-V}$ over $\mathbb{R}^{d}$, $\mathsf{PS}$ considers the augmented distribution $\pi(x,y)\propto\pi^{X}(x)\,\gamma_{h}(y-x)$ over $\mathbb{R}^{d}\times\mathbb{R}^{d}$, where $\gamma_{h}$ denotes the centered Gaussian with covariance $hI_{d}$. Each iteration consists of two steps: for step size $h>0$, (i) $y_{k+1}\sim\pi^{Y|X=x_{k}}=\mathcal{N}(x_{k},hI_{d})$, and (ii) $x_{k+1}\sim\pi^{X|Y=y_{k+1}}$. When $\pi^{X}$ is $\beta$-(log-)smooth, rejection sampling for the backward step $x_{i+1}\sim\pi^{X|Y=y_{i+1}}$ can be performed using $O(1)$ evaluation queries and one proximal query [^1] The proximal query for a function $V:\mathbb{R}^{d}\to\mathbb{R}$ and step size $h>0$ is the function defined as $x\in\mathbb{R}^{d}\mapsto\arg\min\{V(\cdot)+\frac{1}{2h}\,\lVert\cdot-x\rVert^{2}\}$. for any $y_{i+1}$ and $h\asymp(\beta d)^{-1}$.
Even more surprising results were obtained by [CCSW22], which showed that one iteration of $\mathsf{PS}$ can be interpreted as the composition of two continuous-time stochastic processes—Brownian motion and its time-reversal. Since Brownian motion naturally defines a heat semigroup $(P_{t})_{t\geq 0}$ from $L^{2}(\nu*\gamma_{t})$ to $L^{2}(\nu)$, defined by $P_{t}f:=f*\gamma_{t}$, and the time-reversal also admits a (non-homogeneous) semigroup $(Q_{t})_{t\geq 0}$ from $L^{2}(\nu)$ to $L^{2}(\nu*\gamma_{t})$, defined as $Q_{t}f:=P_{t}(f\nu)/P_{t}\nu$ [KP23], this allowed [CCSW22] to carry out a semigroup calculus, leading to contraction results similar to (1.1). For instance, this framework yields a query complexity of $O(qd\beta C_{\mathsf{PI}}(\nu)\,\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu)\log\nicefrac{{1}}{{\varepsilon}})$ for achieving $\varepsilon$-closeness in $\mathcal{R}_{q}$ (i.e., $\mathsf{PS}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ under ($\mathsf{PI}$)). See §3.1 for further details on these semigroups.

### v2

# Zeroth-order Logconcave Sampling

@@ Abstract
We study the zeroth-order query complexity of sampling from a general logconcave distribution: given access to an evaluation oracle for a convex function $V:\mathbb{R}^{d}\rightarrow\mathbb{R}\cup\{\infty\}$, output a point from a distribution within $\varepsilon$-distance to the density proportional to $e^{-V}$. A long line of work provides efficient algorithms for this problem in TV distance, assuming a pointwise warm start (i.e., in $\infty$-Rényi divergence), and using annealing to generate such a warm start. Here, we address the natural and more general problem of using a $q$-Rényi divergence warm start to generate a sample that is $\varepsilon$-close in $q$-Rényi divergence. Our first main result is an algorithm with this end-to-end guarantee with state-of-the-art complexity for $q=\widetilde{\Omega}(1)$. Our second result shows how to generate a $q$-Rényi divergence warm start directly via annealing, by maintaining $q$-Rényi divergence throughout, thereby obtaining a streamlined analysis and improved complexity. Such results were previously known only under the stronger assumptions of smoothness and access to first-order oracles. We also show a lower bound for Gaussian annealing by disproving a geometric conjecture about quadratic tilts of isotropic logconcave distributions.
Central to our approach, we establish hypercontractivity of the heat adjoint and translate this into improved mixing time guarantees for the Proximal Sampler. The resulting analysis of both sampling and annealing follows a simplified and natural path, directly tying convergence rates to isoperimetric constants of the target distribution.
@@ 1 Introduction
We study the zeroth-order query complexity of sampling from an arbitrary logconcave distribution $\pi$ with an arbitrary output guarantee, assuming access to an evaluation oracle for $\log\pi$ up to a fixed unknown additive constant (equivalently, an evaluation oracle for a convex function $V$, with the goal of sampling from the distribution with density proportional to $e^{-V}$). This is a general setting for logconcave sampling—for example, it includes uniform sampling from convex bodies, which will serve as a running example throughout this paper. Results in this general setting provide a theoretical baseline, highlighting the benefits of additional assumptions often made in other sampling work, such as log-smoothness (i.e., $\nabla\log\pi$ is Lipschitz) or access to first-order oracles (i.e., $\nabla\log\pi$). Indeed, much of our work is motivated by gaps between this general case and more favorable settings where first-order oracles are available for unconstrained and log-smooth distributions. In practical terms, zeroth-order samplers are natural default algorithms, since evaluation oracles typically require only function value computations, making them the most general and easiest to implement.
@@
We now make the problem setup more explicit, focusing first on uniform sampling. Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body (i.e., a convex and compact set with nonempty interior) containing the $d$-dimensional unit ball $B_{1}$. Given access to the membership oracle for $\mathcal{K}$ (which answers whether a queried point belongs to $\mathcal{K}$) [GLS93], we ask: how can we sample uniformly from $\mathcal{K}$, and how many oracle queries are required? More formally, given a target accuracy $\varepsilon>0$ and a probability divergence (or distance) $\mathsf{D}$—such as the total variation ($\mathsf{TV}$) distance or $\chi^{2}$-divergence—what is the query complexity of generating a sample whose law is $\varepsilon$-close to the uniform distribution $\pi$ over $\mathcal{K}$ in $\mathsf{D}$?
@@
In this work, we propose a uniform sampler that allows us to establish the first balanced end-to-end complexity guarantee, meaning that the Rényi divergence orders of the initial distribution and the output distribution are the same. This result parallels guarantees found in sampling research in the log-smooth setting. Next, we provide a simple annealing scheme that, for the first time, generates a warm start in $q$-Rényi divergence. To this end, we prove hypercontractivity under simultaneous heat flow, under a logarithmic Sobolev inequality. Finally, we extend these results to arbitrary logconcave distributions under evaluation oracles. Alongside, we present a streamlined analysis leveraging tools from diverse mathematical fields. Combining these contributions, we establish the improved zeroth-order query complexity of sampling from an arbitrary logconcave distribution with $\mathcal{R}_{q}$-guarantees.
1.1 Background and results
@@
In this section, $\pi_{\mathcal{K}}$ denotes the uniform distribution over a convex body $\mathcal{K}$ with diameter $D$, and $\pi$ refers to a logconcave distribution over $\mathbb{R}^{d}$. For clarity, we recall the $q$-Rényi divergence, defined as $\mathcal{R}_{q}(\mu\mathbin{\|}\nu):=\frac{q}{q-1}\log\,\lVert\nicefrac{{\mathrm{d}\mu}}{{\mathrm{d}\nu}}\rVert_{L^{q}(\nu)}$ for $q>1$, and its monotonicity in $q$, and $2\,\lVert\cdot\rVert_{\mathsf{TV}}^{2}\leq\mathsf{KL}=\lim_{q\downarrow 1}\mathcal{R}_{q}\leq\mathcal{R}_{2}=\log(1+\chi^{2})\leq\chi^{2}\leq\mathcal{R}_{\infty}=\log\operatorname{ess\,sup}_{\nu}\nicefrac{{\mathrm{d}\mu}}{{\mathrm{d}\nu}}$. See §1.2 for more details. When a sampler is initialized with a distribution $\nu_{0}$ to sample from a target distribution $\nu$, we define the warmness parameter as $M_{q}=\lVert\mathrm{d}\nu_{0}/\mathrm{d}\nu\rVert_{L^{q}(\nu)}$.
@@ 1.1.1 Uniform sampling from a warm start
The $\mathsf{Ball\ walk}$ and $\mathsf{Hit\text{-}and\text{-}Run}$ are well-studied random walks (i.e., Markov chains) for uniform sampling. These algorithms are analyzed through the conductance framework, where a lower bound on conductance yields a mixing-time guarantee. Establishing such a lower bound involves two ingredients: one-step coupling (i.e., the closeness of transition kernels at two nearby points) and a Cheeger isoperimetric inequality for the target distribution. The Cheeger constant of a probability measure $\pi$ over $\mathbb{R}^{d}$ is defined as
@@
The $\mathsf{Ball\ walk}$ was introduced by Lovász [LOV90]; in each step, it samples a random point uniformly from a ball of fixed radius around the current point, accepting it if the point lies within $\mathcal{K}$. Later, Lovász and Simonovits established that the $\mathsf{Ball\ walk}$ satisfies $\mathcal{R}_{\infty}\to\mathsf{TV}$ with complexity $\operatorname{poly}(d,D,M_{\infty}/\varepsilon)$ [LS93, Remark on p. 398], meaning the initial warmness should be given in $\mathcal{R}_{\infty}$-divergence and the final output is $\varepsilon$-close to $\pi_{\mathcal{K}}$ in $\mathsf{TV}$-distance. They also analyzed an indirect approach that uses the $\mathsf{Ball\ walk}$ (with a Metropolis filter) to sample from a logconcave distribution and then post-processes the sample to bring it closer to $\pi_{\mathcal{K}}$. The resulting algorithm, also of type $\mathcal{R}_{\infty}\to\mathsf{TV}$, was shown to have query complexity $\operatorname{poly}(d,D,\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$.
@@
This approach was further improved by Kannan, Lovász, and Simonovits [KLS97], who introduced the $\mathsf{Speedy\ walk}:\mathcal{R}_{\infty}\to\mathsf{TV}$. Similar to [LS93], it begins by sampling from a biased logconcave variant $\pi^{\prime}$ (of $\pi_{\mathcal{K}}$), which is invariant under the following per-step transition: the next point $x_{i+1}$ is sampled uniformly from $\mathcal{K}\cap B_{\delta}(x_{i})$ via rejection sampling, using the uniform distribution over $B_{\delta}(x_{i})$ as the proposal. Then, a post-processing step is used to reduce the bias. They showed that the mixing time of this chain is $O(d^{2}C_{\mathsf{Ch}}^{-2}(\pi^{\prime})\log\nicefrac{{M_{\infty}}}{{\varepsilon}})$ and that each rejection sampling step requires $O(M_{\infty})$ queries, leading to an overall complexity of $\widetilde{O}(M_{\infty}d^{2}D^{2}\log\nicefrac{{1}}{{\varepsilon}})$ (due to $C_{\mathsf{Ch}}(\pi)\gtrsim D^{-1}$ [LS93]). Hereafter, by following convention, the $\mathsf{Ball\ walk}$ refers to the $\mathsf{Speedy\ walk}$ with the post-processing step. Later, progress on the Kannan-Lovász-Simonovits (KLS) conjecture [KLS95] improved this bound to $C_{\mathsf{Ch}}^{2}(\pi)\gtrsim(\lVert\operatorname{cov}\pi\rVert\log d)^{-1}$ [KLA23]. As a result, the complexity of the $\mathsf{Ball\ walk}$ is now $\widetilde{O}(M_{\infty}d^{2}\lVert\operatorname{cov}\pi\rVert\log\nicefrac{{1}}{{\varepsilon}})$.
@@
The $\mathsf{Hit\text{-}and\text{-}Run}$ algorithm [SMI84] was first analyzed by Lovász [LOV99] via $s$-conductance [LS90], with complexity $\widetilde{O}(M_{\infty}d^{2}D^{2}/\varepsilon^{2})$. Lovász and Vempala [LV06b] showed that the conductance of $\mathsf{Hit\text{-}and\text{-}Run}$ is in fact $\Omega(1/dD)$ resulting in a mixing time of $\widetilde{O}(d^{2}D^{2}\log\nicefrac{{M_{2}}}{{\varepsilon}})$. Each step of $\mathsf{Hit\text{-}and\text{-}Run}$ selects a random chord passing the current point and moves to a random point on the chord. Here, uniform sampling from the chord is implemented via rejection sampling with complexity $O(\log D)$, where the proposal distribution is uniform over a slightly longer chord (found by binary search). They established a novel Cheeger isoperimetric inequality in terms of the cross-cut ratio (rather than Euclidean distance) and showed $\mathsf{Hit\text{-}and\text{-}Run}:\mathcal{R}_{2}\to\mathcal{R}_{2}$ with complexity $\widetilde{O}(d^{2}D^{2}\log\nicefrac{{M_{2}}}{{\varepsilon}})$, making it the first random walk with a balanced guarantee (and to date the only walk with logarithmic dependence on both the initial warmness and the target error). From an $O(1)$-warm start in $\mathcal{R}_{\infty}(\geq\mathcal{R}_{2})$, the $\mathsf{Ball\ walk}$ currently has better complexity than $\mathsf{Hit\text{-}and\text{-}Run}$.
@@
Despite these and many other advances over the past four decades, a unified theory for the complexity of this fundamental problem is missing. First, the $\mathsf{Ball\ walk}$ takes an indirect approach to uniform sampling: it samples from a biased distribution to obtain a mixing guarantee, followed by a post-processing step to reduce the bias. Second, conductance-based analyses require an algorithmic modification making the chains lazy (i.e., the chain remains at its current point with some probability). It would be desirable to develop a sampling algorithm that avoids these additional processing steps. Finally, can we replace the “unbalanced” guarantee of the $\mathsf{Ball\ walk}$ (i.e., $\mathcal{R}_{\infty}\to\mathsf{TV}$) to an arguably more natural $\mathcal{R}_{q}\to\mathcal{R}_{q}$ guarantee for any desired $q$? This question becomes even more natural when considering recent developments in sampling under log-smoothness.
@@
Research in sampling has made substantial progress by analyzing discretizations of continuous-time processes. A central example is the Langevin diffusion $(X_{t})_{t\geq 0}$ in $\mathbb{R}^{d}$, described by the stochastic differential equation
@@
and the smallest such $C$ is called the Poincaré constant $C_{\mathsf{PI}}(\pi)$. The seminal work of Jordan, Kinderlehrer, and Otto [JKO98] interpreted Langevin diffusion as the $2$-Wasserstein gradient flow of $\mathsf{KL}(\cdot\mathbin{\|}\nu)$, revealing a connection to the Logarithmic Sobolev Inequality (LSI), which is stronger than the Poincaré inequality. A probability measure $\pi$ on $\mathbb{R}^{d}$ is said to satisfy a logarithmic Sobolev inequality with constant $C$ if for any locally Lipschitz function $f\in L^{2}(\pi)$,
@@
and the smallest such $C$ is referred to as the log-Sobolev constant $C_{\mathsf{LSI}}(\pi)$. Under ($\mathsf{PI}$) and ($\mathsf{LSI}$) respectively, the Langevin process converges exponentially:
$\chi^{2}(\nu_{t}\mathbin{\|}\nu)\leq\exp\bigl(-\frac{2t}{C_{\mathsf{PI}}(\nu)}\bigr)\,\chi^{2}(\nu_{0}\mathbin{\|}\nu)\,,\quad\text{and}\quad\mathsf{KL}(\nu_{t}\mathbin{\|}\nu)\leq\exp\bigl(-\frac{2t}{C_{\mathsf{LSI}}(\nu)}\bigr)\,\mathsf{KL}(\nu_{0}\mathbin{\|}\nu)\,.$ (1.1)
These bounds were extended to $\mathcal{R}_{q}$-divergence in [CLL19, VW23].
With the assumption of bounded smoothness (i.e., $\beta:=\lVert\nabla^{2}V\rVert<\infty$), a short-time discretization of the Langevin process yields an algorithm that preserves the balanced guarantees of the continuous process. Specifically, it has been shown to converge in $\mathsf{KL}$ (i.e., $\mathcal{R}_{1}\to\mathcal{R}_{1}$) with complexity $\operatorname{poly}(d/\varepsilon,\beta,C_{\mathsf{LSI}}(\nu),\log\mathsf{KL}(\nu_{0}\mathbin{\|}\nu))$ [VW23]. This result was further extended to $\mathcal{R}_{q}$-divergence [CEL+24] achieving a query complexity of $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{PI}}(\nu),\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$ or $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{LSI}}(\nu),\log\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$. We note that the dependence on $\nicefrac{{1}}{{\varepsilon}}$ can be made polylogarithmic by incorporating a Metropolis filter (e.g., $\mathsf{MALA}$ [CLA+21, WSC22, CG23a]) or by considering $\mathsf{MHMC}$ [CV22, CDW+20, CG23b]). For a comprehensive review, we refer readers to the monograph [CHE25].
The complexity of the $\mathsf{Ball\ walk}$ can be described as $\operatorname{poly}(qd/\varepsilon,\beta,C_{\mathsf{PI}}(\nu),\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu))$. In fact, for logconcave probability measures $\pi$, it holds that
@@
The proximal sampler.
Continuing this line of work based on algorithmic diffusion, the $\mathsf{Proximal\ sampler}$ ($\mathsf{PS}$) [LST21] has led to substantial progress on sampling. For a target distribution $\pi^{X}\propto e^{-V}$ over $\mathbb{R}^{d}$, $\mathsf{PS}$ considers the augmented distribution $\pi(x,y)\propto\pi^{X}(x)\,\gamma_{h}(y-x)$ over $\mathbb{R}^{d}\times\mathbb{R}^{d}$, where $\gamma_{h}$ denotes the centered Gaussian with covariance $hI_{d}$. Each iteration consists of two steps: for step size $h>0$,
$\mbox{(i) }y_{k+1}\sim\pi^{Y|X=x_{k}}=\mathcal{N}(x_{k},hI_{d}),\quad\mbox{ and }\quad\mbox{(ii) }x_{k+1}\sim\pi^{X|Y=y_{k+1}}.$
When $\pi^{X}$ is $\beta$-log-smooth, rejection sampling for the backward step $x_{i+1}\sim\pi^{X|Y=y_{i+1}}$ can be performed using $O(1)$ evaluation queries and one proximal query [^1] The proximal query for a function $V:\mathbb{R}^{d}\to\mathbb{R}$ and step size $h>0$ is the function defined as $x\in\mathbb{R}^{d}\mapsto\arg\min\{V(\cdot)+\frac{1}{2h}\,\lVert\cdot-x\rVert^{2}\}$. for any $y_{i+1}$ and $h\asymp(\beta d)^{-1}$.
Further refinements were obtained by [CCS+22], which showed that one iteration of $\mathsf{PS}$ can be interpreted as the composition of two continuous-time stochastic processes—Brownian motion and its time-reversal. Since Brownian motion naturally defines a heat semigroup $(P_{t})_{t\geq 0}$ from $L^{2}(\nu*\gamma_{t})$ to $L^{2}(\nu)$, as $P_{t}f:=f*\gamma_{t}$, and the time-reversal also admits a semigroup $(Q_{t})_{t\geq 0}$ from $L^{2}(\nu)$ to $L^{2}(\nu*\gamma_{t})$, defined as $Q_{t}f:=P_{t}(f\nu)/P_{t}\nu$ [KP23], this allowed [CCS+22] to carry out semigroup calculus, leading to contraction results similar to (1.1). In particular, this framework yields query complexity of $O(qd\beta C_{\mathsf{PI}}(\nu)\,\mathcal{R}_{q}(\nu_{0}\mathbin{\|}\nu)\log\nicefrac{{1}}{{\varepsilon}})$ for achieving $\varepsilon$-closeness in $\mathcal{R}_{q}$ (i.e., $\mathsf{PS}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ under ($\mathsf{PI}$) and smoothness). See §3.1 for further details on these semigroups.